Determine if triangle ABC is a right triangle if it's verticies are A(-9,-2), B(-4,-4), C(-6,-9).
Answer by RAY100(1637)

check slope os AB and BC
m(AB) = ( y2-y1)/(x2-x1)=(-2--4)/(-9--4)=(2)/(-5) =(-2/5)
m(BC) = (-4--9)/(-4--6) = (5/2)
as slopes are negative reciprocals, m(AB)=-1/m(BC), they are perpendicular
as such angle ABC is 90 degrees and triangle is a right triangle by definition